OpenStudy (kikuo):

If you wanted to find how much power you'd need to increase the speed from a physics and mathematical standpoint, how would go about doing so?

OpenStudy (zzr0ck3r):

what?

OpenStudy (kikuo):

Hm, I'm not sure how else to explain. That's how it was explained to me.

OpenStudy (kikuo):

@mathmate

OpenStudy (mathmate):

An object will possess higher kinetic energy at a higher speed than a lower speed. By considering the energy difference, you have an idea of the power/energy required to increase the speed, but this will not account for dissipative losses such as friction and air resistance. These latter will have to be estimated and added on.

OpenStudy (kikuo):

Alright, so since I'm not taking a physics class--this is just for a project--can you explain what kinetic energy is? Also, by what you said, I'm supposed to find the kinetic energy of an object and subtract it from the speed? Sorry, I didn't quite understand haha.

OpenStudy (mathmate):

Ok, here's an example. Kinetic energy of an object of mass m is KE=(1/2)mv^2 v=velocity. A set of consistent units would be m in kg, velocity in m/s. The energy is measured in joules (J), and power in watts (W). So an object of 50 kg moves from velocity 0 to 10 m/s would have initially KE0=(1/2)(50)(0^2)= 0 joules, and finally KE1=(1/2)50(10^2)=2500 joules (unit of energy when using kg and metres) Difference in kinetic energy = KE1-KE0=2500 joules. If this increase in speed (acceleration) is done in 10 seconds, then the power needed is energy difference / time = 2500 joules/10 s. = 250 joules/s = 250 watts = 250 W. (this is the same energy required when you turn on 5 light-bulbs each of 50 W).

OpenStudy (kikuo):

Alright, I'll take this one step at a time. Thanks for sitting and helping me with this lol. What is kinetic energy and velocity exactly?

OpenStudy (mathmate):

Velocity measures how fast a body moves. For energy purposes, you can treat velocity and speed as synonymous. 10 metres per second is like 22 miles per hour. Kinetic energy measures the energy to power up a body to a certain speed/velocity. It is measures in joules. It's the equivalent of gasoline for a car. More gasoline, more miles or acceleration we can give the car. Power is the ability to do work within a given time. For light bulbs, it is measured in watts. For cars, it's measured in horse-powers. Try to read up the following articles, which might help: http://www.physicsclassroom.com/class/energy https://www.youtube.com/watch?v=pDK2p1QbPKQ https://www.khanacademy.org/science/physics/work-and-energy/work-and-energy-tutorial/v/introduction-to-work-and-energy